A financial analyst wished to establish whether the mean return on investment per cent (RoI%)

of the financial companies is greater than the mean RoI% of the manufacturing companies. The

financial analyst randomly sampled 28 financial companies and found their sample mean RoI%

to be 18:714 % with a standard deviation of 9:645%. For a random sample of 24 manufacturing

companies, the sample mean RoI% was 15:125 with a sample standard deviation of 8:823%.

This information is summarized below

Sample 1: Financial Sample 2: Manufacturing

n1 D 28 n2 D 24

x1 D 18:714 x2 D 15:125

S1 D 9:645 S2 D 8:823

Assuming that the population variances are equal, the 95% confidence interval for the difference

between two population means is

1. .5:987I 12:098/

2. .3:589I 5:1838/

3. .0:9678I 4:657/

4. .


1
Expert's answer
2021-08-11T19:33:24-0400

"CI=\bar {X_1}-\bar{X_2}\pm t_{(1-\frac{\alpha}{2}),n_1+n_2-2}sp(\sqrt{\frac{1}{n_1}+\frac{1}{n_2}}"

"sp=\sqrt{\frac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}}"

"\bar X_1=18.714"

"s_1=9.645"

"n_1=28"

"\bar X_2=15.125"

"s_2=8.823"

"n_2=24"

"sp=\sqrt{\frac{(27\times9.645^2)+(23\times 8.823^2)}{50}}"

"=9.276"

"t_{0.025,50}=2.009"

"CI=18.714-15.125\pm(2.009\times 9.276\times\sqrt{\frac{1}{28}+\frac{1}{24}})"

"=3.589\pm5.1827"

=[-1.594,8.772]